# BandRC: Band Shifted Raised Cosine Activated Implicit Neural Representations

### v1

@@ Abstract
In recent years, implicit neural representations(INRs) have gained popularity in the computer vision community. This is mainly due to the strong performance of INRs in many computer vision tasks. These networks can extract a continuous signal representation given a discrete signal representation. In previous studies, it has been repeatedly shown that INR performance has a strong correlation with the activation functions used in its multilayer perceptrons. Although numerous activation functions have been proposed that are competitive with one another, they share some common set of challenges such as spectral bias(Lack of sensitivity to high-frequency content in signals), limited robustness to signal noise and difficulties in simultaneous capturing both local and global features. and furthermore, the requirement for manual parameter tuning. To address these issues, we introduce a novel activation function, Band Shifted Raised Cosine Activated Implicit Neural Networks (BandRC) tailored to enhance signal representation capacity further. We also incorporate deep prior knowledge extracted from the signal to adjust the activation functions through a task-specific model. Through a mathematical analysis and a series of experiments which include image reconstruction (with a +8.93 dB PSNR improvement over the nearest counterpart), denoising (with a +0.46 dB increase in PSNR), super-resolution (with a +1.03 dB improvement over the nearest State-Of-The-Art (SOTA) method for 6X super-resolution), inpainting, and 3D shape reconstruction we demonstrate the dominance of BandRC over existing state of the art activation functions.
@@ 1 Introduction
To address the aforementioned issues, we propose a novel activation function Band Shifted Raised Cosine Activated Implicit Neural Networks (BandRC) which, due to its inherent nature is capable of grasping finer information from signals. In the frequency domain, the proposed activation function takes the form of a frequency-shifted raised cosine filter which offers controlled bandwidth and center frequency while offering frequency compactness. Not only that, but similar to WIRE [19], BandRC offers compactness in spatial domain. Furthermore, we leverage the method introduced by  [9] to obtain the optimal activation function parameters through a task-specific model. Through extensive mathematical and experimental analysis, we show that the BandRC activation function offers superior signal representation and outperforms current state-of-the-art activation functions in terms of higher representation capability and noise robustness. Specifically, we show that our activation function achieves superior performance not only in signal representation but also in other computer vision inverse problems such as denoising, super-resolution, image inpainting.
@@
    Incorporate the leveraging prior knowledge through a task specific model [9] for the automated determination of activation function parameters.
@@ 3.1 BandRC: Implicit Neural Representations with Raised Cosine Activations
We suggest a fixed roll-off rate of $\beta=0.05$ and both bandwidth($T$) and the frequency shift($\zeta$) to be kept as learnable. The outputs at each layer are complex-valued, and the phase is preserved throughout the network. For a stable learning process, we propose to normalize activation function outputs and inputs at each layer and brim them into a unit circle. In the final layer, the real part of the output is extracted. The parameters of each activation function, $T$ and $\zeta$ depend on the training task and should be adjusted accordingly. For this, we leverage some prior knowledge extracted from the training data as suggested in [9]. This prior knowledge embedding process is done through a network block which is composed of a task-specific model followed by a harmonizer network. A latent code $z\in\mathbb{R}^{r}$ is generated from the pre-trained task-specific model and sent to the harmonizer network. The architecture of the harmonizer network was constructed following the approach in  [9], using linear layers followed by SiLU activations.
The harmonizer network then adjusts the activation function parameters in each training iteration. This process eliminates the need for initialization of the activation function parameters, offering a superior benefit over other conventional INR models such as  [19].
shifting parameter ($\zeta$) The frequency shifting component dynamically shifts the base band spectrum of the RC filter, allowing better alignment with high-frequency features allowing better encoding and reconstruction. The prior embeddings feed the $\zeta$ to adhere to the harmonizing atoms to preserve essential frequency components. In other terms the shift repositions signal energy in the Fourier domain. bandwidth parameter ($T$): The bandwidth component adapts the effective bandwidth for each activation. Influencing signal resolution without compromising smoothness and balancing aliasing effects, leading to improved flexibility. Roll-off factor being a constant ensures controlled transition regions, avoiding excessive oscillations. Collectively these independent parameters ensure the capability of the activation to better capture sharp edges, fine textures, and rapid variations in signals, making them more efficient for implicit continuous representations.
@@ 3.2 A mathematical perspective of the activation function
It can be observed that the Raised Cosine function has better stability over a greater number (up to 20) of polynomial coefficients(harmonic integers) comparatively. Both Sinusoidal and Gauss approximations coefficients degrade faster than Raised Cosine. Therefore, reducing the spectral support over a wide range of higher-frequency harmonics.
@@ 4.5 Occupancy Fields
Implicit Neural reconstruction for 3D shapes have been able to encode traditional discrete grid-like coordinate structures to information-rich implicit continuous functions. Thus enabling smooth representations for higher resolutions. The signed distance functions (SDF) provide a negative value for points inside the surface, a positive value for points outside, and zero at the surface. Occupancy fields, on the other hand, provide a binary representation, marking points inside the object as occupied (1) and outside as empty (0). For 3D reconstruction tasks both information forms can be used to train implicit functions.
Data. For this task, we train the Thai statue for SOTA model comparison. First, we create an occupancy volume through point sampling on a 512x512x512 grid, assigning one and zeros to voxels within the object and outside the object respectively.
Architecture. Here, we use resnet3d-18  [28] as the task-specific model truncated up to layer 3. To infuse inherent local features of the whole 3D shape to acquire good the activation parameters appropriately. Thereby handling frequency shifting and filter bandwidth requirements. This enhances better feature representative qualities even in the complex edge-like structures on par with the current state-of-the-art activations.

### v2

@@ Abstract
In recent years, implicit neural representations(INRs) have gained popularity in the computer vision community. This is mainly due to the strong performance of INRs in many computer vision tasks. These networks can extract a continuous signal representation given a discrete signal representation. In previous studies, it has been repeatedly shown that INR performance has a strong correlation with the activation functions used in its multilayer perceptrons. Although numerous activation functions have been proposed that are competitive with one another, they share some common set of challenges such as spectral bias(Lack of sensitivity to high-frequency content in signals), limited robustness to signal noise and difficulties in simultaneous capturing both local and global features. and furthermore, the requirement for manual parameter tuning. To address these issues, we introduce a novel activation function, Band Shifted Raised Cosine Activated Implicit Neural Networks (BandRC) tailored to enhance signal representation capacity further. We also incorporate deep prior knowledge extracted from the signal to adjust the activation functions through a task-specific model. Through a mathematical analysis and a series of experiments which include image reconstruction (with an average PSNR improvement of +5.67 dB over the nearest counterpart across a diverse image dataset), denoising (with a +0.46 dB increase in PSNR), super-resolution (with a +1.03 dB improvement over the nearest State-Of-The-Art (SOTA) method for 6X super-resolution), inpainting, and 3D shape reconstruction we demonstrate the dominance of BandRC over existing state of the art activation functions.
@@ 1 Introduction
To address the aforementioned issues, we propose a novel activation function Band Shifted Raised Cosine Activated Implicit Neural Networks (BandRC) which, due to its inherent nature is capable of grasping finer information from signals. In the frequency domain, the proposed activation function takes the form of a frequency-shifted raised cosine filter which offers controlled bandwidth and center frequency while offering frequency compactness. Not only that, but similar to WIRE [19], BandRC offers compactness in spatial domain. Furthermore, we leverage the method introduced by  [9] to obtain the optimal activation function parameters through a prior-knowledge embedding model. The output of this prior-knowledge model is fed to a special regularization mechanism to boost the optimization process and further enhance the results. Through extensive mathematical and experimental analysis, we show that the BandRC activation function offers superior signal representation and outperforms current state-of-the-art activation functions in terms of higher representation capability and noise robustness. Specifically, we show that our activation function achieves superior performance not only in signal representation but also in other computer vision inverse problems such as denoising, super-resolution, image inpainting.
@@
    Incorporate the leveraging of prior knowledge through a task-specific model [9] for the automated determination of activation function parameters, together with a specialized regularization mechanism.
@@ 3.1 BandRC: Implicit Neural Representations with Raised Cosine Activations
We suggest a fixed roll-off rate of $\beta=0.05$ and both bandwidth($T$) and the frequency shift($\zeta$) to be kept as learnable. The outputs at each layer are complex-valued, and the phase is preserved throughout the network. For a stable learning process, we propose to normalize activation function outputs and inputs at each layer and bring them into a unit circle on the complex plane. In the final layer, the real part of the output is extracted. The parameters of each activation function, $T$ and $\zeta$ depend on the training task and should be adjusted accordingly. For this, we leverage some prior knowledge extracted from the training data as suggested in [9]. This prior knowledge embedding process is done through a network block which is composed of a task-specific model followed by a harmonizer network. A latent code $z\in\mathbb{R}^{r}$ is generated from the pre-trained task-specific model and sent to the harmonizer network. The architecture of the harmonizer network was constructed following the approach in  [9], using linear layers followed by SiLU activations.
The harmonizer network then sends its output to special regularizer blocks attached to each activation function. The regularization mechanism uses a sigmoid projection within some pre-specified bounds 4.
$\theta=a+(b-a).sigmoid(\hat{\theta})$ (4)
Here, a and b are the lower and upper bounds for the activation function parameters($T,\zeta$). They should be pre-specified before training. This harmonizer network, together with the regularization mechanism, then adjusts the activation function parameters in each training iteration. This process eliminates the need for initialization of the activation function parameters, and boosts overall performance, offering a superior benefit over other conventional INR models such as  [19].
The frequency shifting component($\zeta$) dynamically shifts the base band spectrum of the RC filter, allowing better alignment with high-frequency features allowing better encoding and reconstruction. The prior embeddings feed the $\zeta$ to adhere to the harmonizing atoms to preserve essential frequency components. In other terms the shift repositions signal energy in the Fourier domain. The bandwidth component($T$) adapts the effective bandwidth for each activation. Influencing signal resolution without compromising smoothness and balancing aliasing effects, leading to improved flexibility. Roll-off factor($\beta$) being a constant ensures controlled transition regions, avoiding excessive oscillations. Collectively, these independent parameters ensure the capability of the activation to better capture sharp edges, fine textures, and rapid variations in signals, making them more efficient for implicit continuous representations.
@@ 3.2 A mathematical perspective of the activation function
It can be observed that the Raised Cosine function has better stability over a greater number (up to 20) of polynomial coefficients(harmonic integers) comparatively. Both Sinusoidal and Gauss approximated coefficients degrade faster than Raised Cosine. Therefore, reducing the spectral support over a wide range of higher-frequency harmonics.
@@ 3.3 Spectral Bias in INR: Experimental Setup
3.4 Spectral Bias in INR: Layer-wise frequency analysis
A layer-wise frequency analysis was done on BandRC and SIREN activations to experiment with the spectral bias further. A chirp image was used to experiment with the capability of representing high-frequency information as seen in Figure  5. The average magnitude spectrum of each layer output was obtained for the chirp image. The heights of the magnitude spectra along the x-axis of the 2-D Fourier transforms for layers 1-5 are shown in Figure  5. Accordingly, BandRC has higher magnitudes at high frequency components compared to SIREN. As mentioned in  [9], this confirms BandRC is capable in representing better high frequency information.
Figure 5: Layer-wise frequency response for BandRC compared to SIREN
@@ 4.1 Image Reconstruction
Furthermore, we present BandRC’s performance on 24 Kodak lossless images in Fig. 6 (a), which demonstrates it clearly outperforms the SOTA activation functions. It is clear that BandRC consistently achieves the highest performance, averaging 41.24 dB PSNR, significantly exceeding Incode’s 35.57 dB. These results are shown in Table 1. Also, notably, BandRC converges faster compared to other activation functions while maintaining stability as shown in Fig. 6 (b).
Figure 6: BandRC performance analysis on image reconstruction
Activation Average PSNR
ReLU+PE 32.79
SIREN 33.09
WIRE 29.41
INCODE 35.57
BandRC 41.24
Table 1: Average PSNR performance for ImagesBandRC vs. SOTAs in Kodak image reconstruction.
@@ 4.5 Occupancy Fields
Implicit Neural reconstruction for 3D shapes have been able to encode traditional discrete grid-like coordinate structures to information-rich implicit continuous functions. Thus enabling smooth representations for higher resolutions. The signed distance functions (SDF) provide a negative value for points inside the surface, a positive value for points outside, and zero at the surface. Occupancy fields, on the other hand, provide a binary representation, marking points inside the object as occupied (1) and outside as empty (0). For 3D reconstruction tasks, both information forms can be used to train implicit functions.
Data. For this task, we train the Thai statue for SOTA model comparison. First, we create an occupancy volume through point sampling on a 512x512x512 grid, assigning ones and zeros to voxels within the object and outside the object, respectively.
Architecture. Here, we use resnet3d-18  [28] as the task-specific model truncated up to layer 3. To infuse the inherent local features of the whole 3D shape to acquire good activation parameters appropriately. Thereby handling frequency shifting and filter bandwidth requirements. This enhances better feature representative qualities even in the complex edge-like structures on par with the current state-of-the-art activations.
@@
Table 4 reports results for, image representation task with Kodak 22 image, the peak signal-to-noise ratio (PSNR) performance deviations of the proposed BandRC activation for 1000 training epochs under varying architectural configurations. Empirical evidence suggests that the optimal learning rate is sensitive to the model’s depth, and has accordingly been adjusted for each configuration, as indicated in the table. Additionally, we observe a gradual increase in the maximum attainable PSNR with respect to the layer width, highlighting a strong correlation between the number of neurons and the representational capacity induced by the BandRC activation.
Notably, the model achieves a PSNR of approximately 52 dB when configured with a width of 512 neurons and a depth of 4 hidden layers, which is an exceptional performance benchmark in this context. However, to ensure a balance between computational efficiency and reconstruction fidelity, we adopted a configuration with a width of 256 and a depth of 3 hidden layers. This setup yields a PSNR of 39.57 dB, which remains close to the 40 dB threshold and further substantiates the scalability and effectiveness of the proposed nonlinear activation function.
Depth Width 64 128 256 512
2 (lr = 0.01) 28.52 32.92 38.09 42.84
3 (lr = 0.01) 34.40 35.35 39.57 47.80
4 (lr = 0.001) 29.47 35.01 37.19 52.00
Table 4: PSNR ablation study on width/depth in dB